On April!, 2004, a business purchased a machine costing $224 000. The machine has a life expectancy of approximately 40 months. At the end of that time the machine is expected to have a trade-in value of $24 000. The financial year of the business ends on December 31, each year.

(i) Use the formula for calculating annual depreciation on the 'Straight-line Method' to determine the annual depreciation on the machine. ( 2 marks)

(ii) Calculate the depreciation charge at the end of the first year. ( 1 mark)

(iii) Show the depreciation account for the first two years.
Answer #1

a) Annual depreciation = (Original Cost-salvage value)/Useful life = (224000-24000/40) = 5000*12 = 60000

b) Depreciation charge at the end of first year = 60000

c) T account

Depreciation account
Dec 31 60000                
12/31 Bal 60000
Dec 31 60000
12/31 Bal 120000